### v1

# BitNet v2: Native 4-bit Activations with Hadamard Transformation for 1-bit LLMs

Abstract
Efficient deployment of 1-bit Large Language Models (LLMs) is hindered by activation outliers, which complicate quantization to low bit-widths. We introduce BitNet v2, a novel framework enabling native 4-bit activation quantization for 1-bit LLMs. To tackle outliers in attention and feed-forward network activations, we propose $\mathcal{H}$-BitLinear, a module applying an online Hadamard transformation prior to activation quantization. This transformation smooths sharp activation distributions into more Gaussian-like forms, suitable for low-bit representation. Experiments show BitNet v2 trained from scratch with 8-bit activations matches BitNet b1.58 performance. Crucially, BitNet v2 achieves minimal performance degradation when trained with native 4-bit activations, significantly reducing memory footprint and computational cost for batched inference.
There’s Plenty of Room at the Bottom.
Richard Feynman
(a) $\mathbf{W}_{\text{o}}$ of BitNet b1.58
(b) $\mathbf{W}_{\text{down}}$ of BitNet b1.58
(c) $\mathbf{W}_{\text{o}}$ of BitNet v2
(d) $\mathbf{W}_{\text{down}}$ of BitNet v2
Figure 1: Top: Overview of BitNet v2 and $\mathcal{H}$-BitLinear. Bottom: The distribution of the activation of output projection $\mathbf{W}_{\text{o}}$ in attention and down projection $\mathbf{W}_{\text{down}}$ in FFN. BitNet v2 utilizes $\mathcal{H}$-BitLinear to eliminate the large amount of outlier channels in the intermediate states. The Hadamard transformation reshapes the original sharp distribution into a more Gaussian-like form.
1 Introduction
The field of deep learning is rapidly embracing quantization-aware training and low-bit inference, driven by hardware advancements like next-generation GPUs (e.g., GB200) offering native support for 4-bit computations. This promises significant efficiency gains for large-scale models. Pioneering work like BitNet b1.58 [8] demonstrated that 1.58-bit LLMs can match full-precision performance while drastically reducing inference costs (latency, memory, throughput, energy) [16]. However, while BitNet b1.58 quantizes weights to 1.58 bits, alleviating memory bandwidth bottlenecks, it retains 8-bit activations. This reliance on 8-bit precision prevents these models from fully leveraging the 4-bit computational capabilities of emerging hardware, shifting the efficiency bottleneck towards computation itself.
Achieving lower bit-width activations is crucial for maximizing hardware utilization, particularly for efficient kernel design in batched inference scenarios. Research [14, 6] highlights a key challenge: the non-uniform distribution of activations within LLMs. While inputs to attention and feed-forward network (FFN) layers often exhibit Gaussian-like distributions amenable to quantization, their intermediate states (outputs before final projection) contain significant outliers, hindering aggressive low-bit quantization. BitNet a4.8 [14] attempted to address this by selectively using 4-bit quantization for inputs and 8-bit sparsification for intermediate states. While achieving minimal performance loss compared to 8-bit activations, sparsification is less suited for maximizing throughput in batched inference, where dense computations are often preferred for hardware efficiency.
To bridge this gap and unlock the full potential of 4-bit computation for 1.58-bit LLMs, we introduce BitNet v2. Our framework enables native 4-bit activations across the model. The core innovation is $\mathcal{H}$-BitLinear, a novel linear layer replacing the standard output projections in attention and down projections in FFNs. $\mathcal{H}$-BitLinear applies an online Hadamard transformation before activation quantization. This strategically reshapes the sharp, outlier-prone distributions of intermediate states into more manageable, Gaussian-like forms, significantly reducing the impact of outliers in 1.58-bit models. We train BitNet v2 from scratch using 8-bit activations, achieving negligible performance loss compared to BitNet b1.58 [8]. Subsequently, the model can be efficiently fine-tuned with a small amount of data to operate with native 4-bit activations. Extensive experiments demonstrate that our 4-bit BitNet v2 variant achieves performance comparable to BitNet a4.8 while offering superior computational efficiency for batched inference scenarios.
2 BitNet v2: Native 4-bit Activations
We illustrate the architecture of BitNet v2 in Figure 1. We implement BitNet v2 using LLaMA-like components, including RMS normalization [19], SwishGLU [12] and removing all bias. Compared to BitNet [13], we use $\mathcal{H}$-BitLinear for $\mathbf{W}_{\text{o}}$ in attention and $\mathbf{W}_{\text{down}}$ in FFN layers to deal with outlier channels of intermediate states. BitNet v2 is trained with 1.58-bit weights and INT8 activations from scratch, then continue-trained with INT4 activations for all linear layers except input/output embedding.
2.1 $\mathcal{H}$-BitLinear
Following [13, 8], as for weight quantization, we use per-tensor absmean function to quantize the weights into ternary values, i.e., {-1, 0, 1}:
$\text{Q}_{w}(\mathbf{W})=\alpha\text{RoundClip}(\frac{\mathbf{W}}{\alpha+\epsilon},-1,1),\,\alpha=\text{mean}(|\mathbf{W}|)$ (1)
$\text{RoundClip}(X,a,b)=\min(\max(\text{round}(X),a),b)$ (2)
For low-bit activations, previous works [14, 6] have shown that the distributions of the inputs to the attention and feed-forward-network layers (i.e., the activations of $\mathbf{W}_{\text{qkv}}$ and $\mathbf{W}_{\text{up,gate}}$) tend to exhibit a Gaussian-like shape, while the intermediates states (i.e., the activations of $\mathbf{W}_{\text{o}}$ and $\mathbf{W}_{\text{down}}$) have more outlier channels and massive amount of entries around zero.
Therefore, we introduce $\mathcal{H}$-BitLinear for $\mathbf{W}_{\text{o}}$ in attention and $\mathbf{W}_{\text{down}}$ in FFN layers. $\mathcal{H}$-BitLinear employs a Hadamard transformation before activation quantization to first reduce the number of outlier channels. The Hadamard transformation satisfies that:
$\text{Hadamard}(\mathbf{X})=\mathbf{H_{m}}\mathbf{X}$ (3)
$\mathbf{H_{m}}=\cfrac{1}{\sqrt{2}}\left(\begin{array}[]{cc}\mathbf{H_{m-1}}&\mathbf{H_{m-1}}\\ \mathbf{H_{m-1}}&-\mathbf{H_{m-1}}\end{array}\right),\,\mathbf{H_{0}}=(1)$ (6)
where $\mathbf{H_{m}}$ is a $2^{m}\times 2^{m}$ matrix and $\mathbf{X}\in\mathbf{R}^{n}$, $n=2^{m}$. We use fast-hadamard-transform [^1] https://github.com/Dao-AILab/fast-hadamard-transform to perform the matrix multiplication, which has $\mathcal{O}(n\log n)$ computational complexity.
As shown in Figure 2 and Figure 3, with Hadamard transformation, the distribution of the intermediate states becomes closer to a Gaussian-like distribution, which significantly reduces the number of outliers and make it more suitable for INT4 quantization.
For 8-bit and 4-bit activations, we adopt per-token absmax and absmean function, respectively. The activation quantization can be formulated as:
$\text{Q}_{\text{INT8}}(\mathbf{X})=\frac{\gamma}{127}\text{RoundClip}(\frac{127}{\gamma+\epsilon}\mathbf{X},-128,127),\,\gamma=\max(|\mathbf{X}|)$ (7)
$\text{Q}_{\text{INT4}}(\mathbf{X})=\frac{\beta}{\sqrt{7}}\text{RoundClip}(\frac{\sqrt{7}}{\beta+\epsilon}\mathbf{X},-8,7),\,\beta=\text{mean}(|\mathbf{X}|)$ (8)
Above all, the matrix multiplication of $\mathcal{H}$-BitLinear can be written as:
$\mathbf{Y}=\text{Q}_{w}(\mathbf{W})\cdot\text{Q}_{\text{INT8/4}}(\mathbf{X_{r}}),\,\,\mathbf{X_{r}}=\text{Hadamard}(\text{LN}(\mathbf{X}))$ (9)
where LN denotes the layer normalization.
(a) $\mathbf{W}_{\text{qkv}}$ of BitNet b1.58
(b) $\mathbf{W}_{\text{o}}$ of BitNet b1.58
(c) $\mathbf{W}_{\text{up,gate}}$ of BitNet b1.58
(d) $\mathbf{W}_{\text{down}}$ of BitNet b1.58
(e) $\mathbf{W}_{\text{qkv}}$ of BitNet v2
(f) $\mathbf{W}_{\text{o}}$ of BitNet v2
(g) $\mathbf{W}_{\text{up,gate}}$ of BitNet v2
(h) $\mathbf{W}_{\text{down}}$ of BitNet v2
Figure 2: The activation distribution of BitNet b1.58 and BitNet v2 with 8-bit activations.
(a) $\mathbf{W}_{\text{down}}$ of BitNet b1.58
(b) $\mathbf{W}_{\text{down}}$ of BitNet v2
(c) $\mathbf{W}_{\text{o}}$ of BitNet b1.58
(d) $\mathbf{W}_{\text{o}}$ of BitNet v2
Figure 3: The activation distribution of $\mathbf{W}_{\text{down}}$ in FFN and $\mathbf{W}_{\text{o}}$ in attention of BitNet b1.58 and BitNet v2 with 8-bit activations.
[…]
BitNet b1.58 400M 13.37 24.32 43.01 39.51 64.91 51.93 45.51 44.87
BitNet a4.8 13.61 24.15 41.75 39.48 65.18 53.59 44.34 44.75
BitNet v2 (a8) 13.50 23.29 43.06 39.06 64.74 50.59 45.26 44.33
BitNet v2 (a4) 13.78 23.29 41.46 38.33 65.45 50.59 44.56 43.95
BitNet b1.58 1.3B 11.02 27.90 49.58 48.85 69.80 55.80 54.12 51.01
BitNet a4.8 11.15 27.47 49.20 48.72 69.64 56.51 53.85 50.90
BitNet v2 (a8) 11.14 27.90 49.96 48.37 69.42 57.22 54.14 51.17
BitNet v2 (a4) 11.33 27.56 49.58 48.00 68.23 55.49 53.58 50.41
BitNet b1.58 3B 9.71 28.84 54.80 56.39 71.44 59.35 60.47 55.22
BitNet a4.8 9.80 29.01 55.01 55.92 71.76 59.59 59.85 55.19
BitNet v2 (a8) 9.72 30.55 55.56 57.19 71.33 58.72 60.90 55.71
BitNet v2 (a4) 9.85 28.92 55.01 56.59 71.65 59.67 60.74 55.43
BitNet b1.58 7B 9.09 31.74 59.51 61.49 74.37 59.98 61.63 58.12
BitNet a4.8 9.16 31.91 59.09 61.06 74.16 59.67 61.54 57.91
BitNet v2 (a8) 9.14 32.94 58.54 61.08 74.10 61.48 64.22 58.73
BitNet v2 (a4) 9.24 32.42 58.00 60.71 74.27 60.85 63.52 58.30
Table 1: Perplexity and results of BitNet v2, BitNet a4.8 and BitNet b1.58 on the end tasks.
2.2 Training
Following [13], we employ the straight-through estimator (STE) [2] for gradient approximation and use mixed-precision training to update the parameters. During backward propagation, we bypass the non-differentiable functions in quantization. To support mixed-precision training, we maintain a full-precision latent weight to accumulate parameter updates.
For the backward pass through the Hadamard transformation, we apply the transformation to the gradients as well, leveraging the orthogonality of the transformation matrix $\mathbf{H_{m}}$. Specifically, the backward propagation is formulated as:
$\cfrac{\partial\mathcal{L}}{\partial\mathbf{X}}=\text{Hadamard}(\cfrac{\partial\mathcal{L}}{\partial\,\text{Hadamard}(\mathbf{X})})$ (10)
Similar to BitNet a4.8, BitNet v2 with 4-bit activations can be continue-trained from its 8-bit activation counterpart using a small number of training tokens, while incurring negligible performance loss. The optimizer states are reused for continue-training.
3 Experiments
We compared BitNet v2 to BitNet b1.58 and BitNet a4.8 of various model sizes. All models were trained with 1.58-bit weights. BitNet b1.58 has fully INT8 activations for all linear layers. BitNet a4.8 is continue-trained from BitNet b1.58 with a hybrid quantization and sparsification for activations, where the inputs to the sublayers are quantized into 4-bit integers and the intermediate states uses top-$K$ sparsification [15] and squared ReLU.
We adopted the two-stage weight decay and learning rate scheduling following the training recipe of BitNet b1.58 [8]. All models were trained with 100B tokens from the RedPajama dataset [4] to ensure a fair comparison. For BitNet v2 (a4) and BitNet a4.8, we first trained the model with 8-bit activations for 95B tokens. Then we reused the optimizer states and continue-train the model with 4-bit activations for 5B tokens. More details can be found in the Appendix B.
We evaluated the zero-shot accuracy for these models on a range of language tasks using the lm-evaluation-harness toolkit [5], including ARC-Easy (ARCe) [17], ARC-Challenge (ARCc) [17], Hellaswag (HS) [18], Winogrande (WGe) [11] and PIQA (PQ) [3] and LAMBADA (LBA) [9]. We also reported the perplexity on the validation set of C4 [10] dataset.
[…]
BitNet v2 (a8) 3B 30.55 55.56 57.19 71.33 58.72 60.90 55.71
w/ 4-bit KV 29.52 55.18 57.17 70.95 58.56 60.84 55.37
w/ 4-bit QKV 30.63 55.22 57.15 71.16 58.96 60.49 55.60
w/ 4-bit Q, 3-bit KV 29.69 55.22 56.22 71.49 57.62 59.01 54.88
BitNet v2 (a8) 7B 32.94 58.54 61.08 74.10 61.48 64.22 58.73
w/ 4-bit KV 33.02 58.67 61.04 73.61 61.88 64.06 58.71
w/ 4-bit QKV 32.76 58.46 61.01 74.10 60.85 63.87 58.51
w/ 4-bit Q, 3-bit KV 32.51 58.29 60.85 73.39 60.77 62.99 58.13
Table 2: The zero-shot accuracy of BitNet v2 with 8-bit activations and QKV states varying bit-widths on the end tasks.
[…]
BitNet v2 (a4) 3B 28.92 55.01 56.59 71.65 59.67 60.74 55.43
w/ 4-bit KV 29.52 54.46 56.36 71.49 58.17 60.14 55.02
w/ 4-bit QKV 28.58 55.43 56.32 71.16 57.93 60.70 55.02
w/ 4-bit Q, 3-bit KV 29.18 55.51 55.85 71.60 58.41 59.54 55.02
BitNet v2 (a4) 7B 32.42 58.00 60.71 74.27 60.85 63.52 58.30
w/ 4-bit KV 32.94 58.12 60.33 74.21 61.01 63.65 58.38
w/ 4-bit QKV 33.11 57.91 60.78 74.05 61.17 62.93 58.33
w/ 4-bit Q, 3-bit KV 32.08 57.95 60.29 73.23 59.59 62.97 57.69
Table 3: The zero-shot accuracy of BitNet v2 with 4-bit activations and QKV states varying bit-widths on the end tasks.
3.1 Main Results
We present the detailed results of BitNet v2 and the baselines in Table 1. Introducing the Hadamard transformation before the quantization in attention and FFN layers results in minimal perplexity degradation. For 8-bit activations, BitNet v2 surpasses BitNet b1.58 with an average accuracy improvement of 0.16%, 0.49%, and 0.61% on end tasks for the 1.3B, 3B, and 7B model sizes, respectively. Additionally, BitNet v2 enables native 4-bit activations across all linear layers, enhancing efficiency for batched inference. With INT4 activations, BitNet v2 achieves perplexity comparable to BitNet a4.8 while demonstrating superior performance on downstream tasks for the 3B and 7B models.
Table 2 and Table 3 summarize detailed results of BitNet v2 (a8) and BitNet v2 (a4) with low-bit attention, respectively. We adopt post-RoPE quantization for QKV states. The QKV heads were directly quantized to unsigned integers using the absmax function, without the need of any calibration dataset. We retain the KV heads of [BOS] token as 8-bit precision. As shown in Table 2 and Table 3, BitNet v2 with 3-bit KV Cache achieves accuracy comparable to its counterpart with full-precision KV cache in 3B and 7B models.
3.2 Comparison with Post-Training Quantization.
We compared BitNet v2 (a4) with post-training quantization baselines, including SpinQuant [7] and QuaRot [1], in 1.3B models. QuaRot employs randomized Hadamard transformations to mitigate outlier features, while SpinQuant uses the learnable rotary matrix. Then they adopt GPTQ and absmax function to quantize the weight and activations into 4-bit, respectively. Since the weights of BitNet b1.58 were already trained to be ternary values from scratch, we adopted absmean function used in the training of BitNet rather than GPTQ for weight quantization. For activation quantization of the baselines, we retained rotary transformations for activations as an online operation; Furthermore, we removed the fusion of RMS normalization scales into projections. Following [7], we tuned the rotation of SpinQuant with 800 samples from WikiText2 dataset.
We report the results of BitNet v2 (a4) and BitNet b1.58 with SpinQuant and QuaRot in Table 4. All models were quantized to 1.58-bit weight and 4-bit activations. BitNet v2 (a4) significantly surpasses these baselines in terms of perplexity on valid set of C4 and accuracy on downstream tasks. Additionally, we observed that ternary models are more sensitive to the fusion of rotary matrix and latent weights. Specifically, we removed rotary matrix fusion for $\text{W}_{\text{qkv}}$ in attention modules and $\text{W}_{\text{up, gate}}$ in FFNs, i.e., using $\text{Q}_{\text{w}}(W)R^{T}$ rather than $\text{Q}_{\text{w}}(WR^{T})$. Removing this fusion notably enhances baseline performance: the perplexity decreases from 19.80 to 13.52 for SpinQuant. However, they still trails substantially behind BitNet v2 (a4). Moreover, this adjustment forces these projections to revert to full precision (W16A4), thus sacrificing inference efficiency.
[…]
w/o fusing rotary matrix to $\text{W}_{\text{qkv,up,gate}}$
QuaRot 13.52 26.28 47.43 45.92 65.89 51.46 42.34 46.55
SpinQuant 13.52 25.60 47.35 45.52 67.25 52.49 42.52 46.79
QuaRot 20.83 24.74 40.78 40.54 62.89 49.33 36.89 42.53
SpinQuant 19.80 24.74 40.19 40.77 62.73 52.09 39.24 43.29
BitNet v2 (a4) 11.33 27.56 49.58 48.00 68.23 55.49 53.58 50.41
Table 4: Perplexity and zero-shot accuracy of BitNet v2, QuaRot and SpinQuant on the end tasks.
3.3 Ablation Study
We conduct ablation studies on the Hadamard transformation in $\mathcal{H}$-BitLinear at the 1.3B and 3B model scales. All models are trained on the same dataset to ensure a fair comparison. We report perplexity on the C4 validation set and zero-shot accuracy across a range of language tasks, including ARC-Easy [17], ARC-Challenge [17], Hellaswag [18], Winogrande [11], PIQA [3], and LAMBADA [9]. As shown in Table 5, removing the rotary transformation leads to model divergence. Moreover, while applying the Hadamard transformation to both the weights and activations results in faster convergence, it achieves similar performance to applying it only to the activations as training progresses. Therefore, for simplicity, we apply the Hadamard transformation only to the activations in $\mathcal{H}$-BitLinear. Detailed results can be found in Appendix A.
Methods #Bits 1.3B 3B
[…]
No rotatation W1.58A8 diverged diverged
Weight & activation rotatation 50.47 11.14 55.55 9.69
Activation rotatation 51.16 11.14 55.71 9.72
No rotatation W1.58A4 diverged diverged
Weight & activation rotatation 50.09 11.33 54.98 9.81
Activation rotatation 50.41 11.33 55.43 9.85
Table 5: Ablations on the Hadamard transformation of $\mathcal{H}$-BitLinear across various sizes.
4 Conclusion
We introduce BitNet v2, enabling native 4-bit activations within 1-bit LLMs. This is achieved using our proposed $\mathcal{H}$-BitLinear layer in place of standard attention output and FFN down projections. $\mathcal{H}$-BitLinear employs an online Hadamard transformation before activation quantization to effectively suppress outlier channels by reshaping the activation distribution. Our experiments show BitNet v2 with 8-bit activations matches BitNet b1.58 performance. Subsequently, BitNet v2 can be trained for native 4-bit activation use. This 4-bit variant, BitNet v2(a4), maintains comparable performance to the 8-bit version while significantly boosting efficiency in batched inference scenarios.
Appendix A More results
[…]
No rotatation 1.3B diverged
Weight & activation rotatation 27.13 50.29 48.32 69.04 54.30 53.76 50.47
Activation rotatation 27.90 49.96 48.37 69.42 57.22 54.14 51.17
No rotatation 3B diverged
Weight & activation rotatation 30.03 55.72 56.81 71.65 59.43 59.65 55.54
Activation rotatation 30.55 55.56 57.19 71.33 58.72 60.90 55.71
Table 6: Ablations on the Hadamard transformation of $\mathcal{H}$-BitLinear across various sizes. All models have 1.58-bit weights and 8-bit activations.
[…]
No rotatation 1.3B diverged
Weight & activation rotatation 27.22 49.12 47.77 69.37 54.54 52.49 50.09
Activation rotatation 27.56 49.58 48.00 68.23 55.49 53.58 50.41
No rotatation 3B diverged
Weight & activation rotatation 29.44 54.46 56.57 71.93 57.85 59.64 54.98
Activation rotatation 28.92 55.01 56.59 71.65 59.67 60.74 55.43
Table 7: Ablations on the Hadamard transformation of $\mathcal{H}$-BitLinear across various sizes. All models have 1.58-bit weights and 4-bit activations.
Appendix B Hyper-parameters
Size Hidden Size GLU Size #Heads #Layers Batch Size # Tokens Seq Length
400M 1024 4096 16 24 1M 100B 2048
1.3B 2048 8192 32 18 1M 100B 2048
3B 4096 8192 32 20 1M 100B 2048
7B 4096 16384 32 24 1M 100B 2048
Table 8: Model configurations for the BitNet models.
Model Size Learning Rate Weight Decay Warm-up Adam $\beta$
BitNet 400M $1.8\times 10^{-3}\rightarrow 1.2\times 10^{-3}$ $0.1\rightarrow 0$ 375 (0.9, 0.95)
1.3B $1.2\times 10^{-3}\rightarrow 8\times 10^{-4}$ $0.1\rightarrow 0$ 375 (0.9, 0.95)
3B $1.2\times 10^{-3}\rightarrow 6.4\times 10^{-4}$ $0.1\rightarrow 0$ 375 (0.9, 0.95)
7B $1\times 10^{-3}\rightarrow 6\times 10^{-4}$ $0.1\rightarrow 0$ 375 (0.9, 0.95)
Table 9: Hyper-parameters for both BitNet v2 training.

### v2

@@
 […]
-No rotatation W1.58A8 diverged diverged
-Weight & activation rotatation 50.47 11.14 55.55 9.69
-Activation rotatation 51.16 11.14 55.71 9.72
-No rotatation W1.58A4 diverged diverged
-Weight & activation rotatation 50.09 11.33 54.98 9.81
-Activation rotatation 50.41 11.33 55.43 9.85
+No rotation W1.58A8 51.01 11.02 55.22 9.71
+Weight & activation rotation 50.47 11.14 55.55 9.69
+Activation rotation 51.16 11.14 55.71 9.72
+No rotation W1.58A4 diverged diverged
+Weight & activation rotation 50.09 11.33 54.98 9.81
+Activation rotation 50.41 11.33 55.43 9.85
 Table 5: Ablations on the Hadamard transformation of $\mathcal{H}$-BitLinear across various sizes.
@@
 […]
-No rotatation 1.3B diverged
-Weight & activation rotatation 27.13 50.29 48.32 69.04 54.30 53.76 50.47
-Activation rotatation 27.90 49.96 48.37 69.42 57.22 54.14 51.17
-No rotatation 3B diverged
-Weight & activation rotatation 30.03 55.72 56.81 71.65 59.43 59.65 55.54
-Activation rotatation 30.55 55.56 57.19 71.33 58.72 60.90 55.71
+No rotation 1.3B diverged
+Weight & activation rotation 27.13 50.29 48.32 69.04 54.30 53.76 50.47
+Activation rotation 27.90 49.96 48.37 69.42 57.22 54.14 51.17
+No rotation 3B diverged
+Weight & activation rotation 30.03 55.72 56.81 71.65 59.43 59.65 55.54
+Activation rotation 30.55 55.56 57.19 71.33 58.72 60.90 55.71
 Table 6: Ablations on the Hadamard transformation of $\mathcal{H}$-BitLinear across various sizes. All models have 1.58-bit weights and 8-bit activations.
 […]
-No rotatation 1.3B diverged
-Weight & activation rotatation 27.22 49.12 47.77 69.37 54.54 52.49 50.09
-Activation rotatation 27.56 49.58 48.00 68.23 55.49 53.58 50.41
-No rotatation 3B diverged
-Weight & activation rotatation 29.44 54.46 56.57 71.93 57.85 59.64 54.98
-Activation rotatation 28.92 55.01 56.59 71.65 59.67 60.74 55.43
+No rotation 1.3B diverged
+Weight & activation rotation 27.22 49.12 47.77 69.37 54.54 52.49 50.09
+Activation rotation 27.56 49.58 48.00 68.23 55.49 53.58 50.41
+No rotation 3B diverged
+Weight & activation rotation 29.44 54.46 56.57 71.93 57.85 59.64 54.98
+Activation rotation 28.92 55.01 56.59 71.65 59.67 60.74 55.43
 Table 7: Ablations on the Hadamard transformation of $\mathcal{H}$-BitLinear across various sizes. All models have 1.58-bit weights and 4-bit activations.